OpenStudy (bbb911):

This is Geometry: Suppose a chord in a circle is 80 centimeters long and it is 30 centimeters from the center of the circle. Find the measure of a radius of the circle.

OpenStudy (anonymous):

50 cm

OpenStudy (bbb911):

wow awesome looks like you didn't have to think about it thanks so much :D

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

r=30^2+40^2= r= 50

OpenStudy (anonymous):

80/2=40

OpenStudy (bbb911):

:)

OpenStudy (anonymous):

ty

OpenStudy (bbb911):

ur welcome